A plane flies in a loop at 1000km/h. At the top of the loop the pilot feels a force from their chair which is double their weight. What is the radius of the loop?

a) 1.593km

b) 2.625km

c) 2.928km
Answer #1

From Newton's second law

FN = mV2/r - mg

Given FN =2 mg

2mg = mV2/r - mg

3mg = mV2/r

r = V2/3g

Given V=1000*1000/3600 =277.78 m/s

r=277.782/3*9.8

r = 2.625 km